OpenStudy (anonymous):

Classify the system of equations. -1/3x+y+2=0, 2x+y+5=0 (Would it be parallel, coincident, or intersecting?) @bibby @StudyGurl14 @Abbot @dan815

OpenStudy (anonymous):

@geekfromthefutur

OpenStudy (geekfromthefutur):

Do you have a guess what it might be?

OpenStudy (anonymous):

Coincident maybe,, but I honestly I don't know how to do this

OpenStudy (bibby):

get them into point intercept form -1/3x+y+2=0, 2x+y+5=0 \(y=\frac{1}{3}x-2\) \(y=-2x-5\) they're definitely not parallel

OpenStudy (geekfromthefutur):

I agree with @bibby its not parallel and its not intersecting because there isn't two lines. I know that because there isnt two equations as well so wat does that nail it down to?

OpenStudy (bibby):

there are 2 lines. they aren't coincidental because they aren't the same lines

OpenStudy (bibby):

at least according to what google says coincidental lines are

OpenStudy (geekfromthefutur):

@Blu-Girl

OpenStudy (anonymous):

hmm clicked coincidental wasn't it, but I don't blame y'all cuz I guessed the same thing

OpenStudy (bibby):

it was intersecting

OpenStudy (bibby):

I didn't guess rofl

OpenStudy (anonymous):

oh lol

OpenStudy (bibby):

a) they have different slopes b) they have different equations c) they're not parallel

OpenStudy (anonymous):

oohh

OpenStudy (geekfromthefutur):

This might help

OpenStudy (geekfromthefutur):

thats the problem on a graph

OpenStudy (anonymous):

I'll look at it

OpenStudy (bibby):

so put the 2 thoughts together parallel, coincident, or intersecting a) they have different slopes, they aren't parallel b) they have different equations, they aren't coincident c) they're not parallel, 2 lines with different slopes will eventually intersect

OpenStudy (anonymous):

Thks

OpenStudy (geekfromthefutur):

I think its consident

OpenStudy (bibby):

We've come to the conclusion it wasn't. it's intersecting.

OpenStudy (anonymous):

yea.... I know bibby...

OpenStudy (anonymous):

:)

OpenStudy (geekfromthefutur):

o ok

OpenStudy (bibby):

:)